A manager of a candy shop sells chocolate peanuts for $9 per pound and chocolate covered cashews for $15 per pound. The manager wishes to mix 50 pounds of cashews to get a cashew-peanut mixture that sell for 9.50 per pound. How many pounds of chocolate covered peanuts should be used?
Answer by jorel1380(3719)

50(15)+m(9)=9.50(50+m)
750+9m=475+9.5m
275=.5m
m=275/.5=550
The manager needs 550 lbs. of chocolate covered peanuts for his mix..